Grade 12Algebra

Find
0.35 ÷ 5
Find
0.35 ÷ 5
Find
0.35 ÷ 5
Find
0.35 ÷ 5

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (35/100) ÷ 5
Then,
= (35/100) × (1/5)
= (7/100) × (1/1)
= (7/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places.
Then,
= 0.07

Thanks